Factor the following polynomial. Reduce any fractions.
X^3-2x^2-5x+6

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Factor the following polynomial. Reduce any fractions.
X^3-2x^2-5x+6
===============
What would you learn from me doing it?

Answer by ikleyn(52781)

.
x^3 - 2x^2 - 5x + 6 = x^3 - 3x^2 + x^2 - 5x + 6 = (x^3-3x^2) + (x^2 - 5x + 6) = x^2*(x-3) + (x-2)*(x-3) = (x-3)*(x^2 + (x-2)) =

= (x-3)*(x^2+x-2) = (x-3)*(x+2)*(x-1).

Solved.